# Tail Distribution of Regret in Optimistic Reinforcement Learning

### v2

@@ 5.1 Proof of Theorem 1
Proof of Theorem 1.
Recall that $\bm{\pi}^{*}=\{\pi^{*}_{h}\}_{h=0}^{H-1}\in{\mathbf{\Pi}}^{*}$, where $\mathbf{\Pi}$$^{*}$ denotes the set of optimal policies for the transition kernel $\bm{P}^{*}=\{P^{*}_{h}\}_{h=0}^{H-2}$. At episode $k$, the algorithm computes a greedy policy $\pi^{k}=\{\pi^{k}_{h}\}_{h=0}^{H-1}$ using an optimistic estimate of the optimum value function $\{V^{*}_{h}\}_{h=0}^{H-1}$ based on data collected up to episode $k$. Recall that $R_{K}$ denote the cumulative regret incurred up to episode K1K-1. Then, writing $s_{0}^{k}\equiv s_{0}$, we have:
@@
Note that for any non negative numbers $a_{1},a_{2},\dots,a_{H}$, the $\mathrm{clip}$ function satisfies
$\mathrm{clip}\left(\sum_{i=1}^{H}a_{i}\right)\leq\sum_{i=1}^{H}\mathrm{clip}(Ha_{i}).$ (4)
@@
$=\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})+\Delta^{k}_{h+1}(s_{h+1}^{k})-\eta^{k}_{h+1}(s_{h+1}^{k}),$
where the sequence $\{\eta^{k}_{h}(s_{h}^{k})\}_{h=1,k=0}^{H-1,K-1}$ defined as
$\eta^{k}_{h}(s_{h}^{k})=\Delta^{k}_{h}(s_{h}^{k})-\mathbb{E}\left[\Delta^{k}_{h}(s_{h}^{k})\big|\mathcal{F}_{h-1}^{k}\right]$
is a Martingale difference sequence with the filtration $\{\mathcal{F}_{h-1}^{k}\}_{h=1,k=0}^{H-1,K-1}$. For the sake of notational simplicity, we define $\eta^{k}_{H}\equiv 0$ for all $k\in[K]$. Using a telescopic sum, noting that $\Delta^{k}_{H}\equiv 0$, we obtain
$\Delta^{k}_{0}(s_{0}^{k})$ $=\sum_{h=0}^{H-1}\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})$
$=\sum_{h=0}^{H-1}\mathrm{clip}\left(\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})\right)-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})$ (5)
where the last equality follows by the fact that $\mathrm{gap}_{h}(s,a)\geq\mathrm{gap}^{*}$ if $a\notin\mathcal{A}^{*}_{h}(s)$ and $\mathrm{gap}_{h}(s,a)=0$ otherwise.
Note that $\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})=V_{h}^{*}(s_{h}^{k})-Q_{h}^{*}(s_{h}^{k},a_{h}^{k})\leq V_{h}^{*}(s_{h}^{k})-V_{h}^{k}(s_{h}^{k})+V_{h}^{k}(s_{h}^{k})-V_{h}^{\bm{\pi}^{k}}(s_{h}^{k})$. The bonus function $b_{h}^{k}(\cdot)$ is chosen so that $V_{h}^{k}$ provides an optimistic estimate of $V_{h}^{*}$, as long as the estimated transition kernel $\hat{P}_{h^{\prime}}^{k}$ and the true transition kernel $P_{h^{\prime}}^{*}$ remain close for all horizons $h^{\prime}\geq h$. This combined with the monotonicity of the $\mathrm{clip}(\cdot)$ function provides an upper bound for the regret at episode $k\geq\lceil K^{\gamma}\rceil$. We define the “good event at episode kk” as
@@
We proceed with bounding the accumulative regret $R_{K}^{(1)}=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left(V_{0}^{*}(s_{0}^{k})-V_{0}^{\bm{\pi}_{k}}(s_{0}^{k})\right)$, under the event $\bar{\mathcal{G}}^{K}$.
$R_{K}^{(1)}\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$
$=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left[\sum_{h=0}^{H-1}\mathrm{clip}\left(\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})\right)-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ (by (5))
$=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left[\sum_{h=0}^{H-1}\mathrm{clip}\left(\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})\right)\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$
$=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\Bigg[\sum_{h=0}^{H-1}\mathrm{clip}\left(V_{h}^{*}(s_{h}^{k})-V_{h}^{k}(s_{h}^{k})+V_{h}^{k}(s_{h}^{k})-Q_{h}^{*}(s_{h}^{k},a_{h}^{k})\right)\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}\mathbbm{1}\{\mathcal{G}^{k}\}$
$\qquad\qquad\qquad\qquad-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\Bigg]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$
$\leq\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\Bigg[\sum_{h=0}^{H-1}\mathrm{clip}\left(V_{h}^{*}(s_{h}^{k})-V_{h}^{k}(s_{h}^{k})+V_{h}^{k}(s_{h}^{k})-V_{h}^{\bm{\pi}^{k}}(s_{h}^{k})\right)\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}\mathbbm{1}\{\mathcal{G}^{k}\}$
$\qquad\qquad\qquad\qquad-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\Bigg]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ ($V_{h}^{\bm{\pi}^{k}}(s_{h}^{k})=Q_{h}^{\bm{\pi}^{k}}(s_{h}^{k},a_{h}^{k})\leq Q_{h}^{*}(s_{h}^{k},a_{h}^{k})$)
$\leq\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left[\sum_{h=0}^{H-1}\mathrm{clip}\left(V_{h}^{k}(s_{h}^{k})-V_{h}^{\bm{\pi}^{k}}(s_{h}^{k})\right)\mathbbm{1}\{\mathcal{G}^{k}\}\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ (by (6))
$=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left[\sum_{h=0}^{H-1}\mathrm{clip}\left(\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}\right)\mathbbm{1}\{a_{h}^{k}\notin\mathcal{A}_{h}^{*}(s_{h}^{k})\}-\sum_{h=0}^{H-1}\eta^{k}_{h+1}(s_{h+1}^{k})\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\},$ (7)
@@
$=2\mathbbm{1}\{\mathcal{G}^{k}\}b^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))+\tilde{\Delta}_{h+1}^{k}(s_{h+1}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}-\tilde{\eta}_{h+1}^{k}(s_{h+1}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}.$ (8)
@@
where $(a)$ follows from Hölder’s inequality, and noting that $\sum_{s}(P_{h}^{*}(s\mid s_{h}^{k},a_{h}^{k})-\hat{P}_{h}^{k}(s\mid s_{h}^{k},a_{h}^{k}))=0$, and the sequence $\{\tilde{\eta}^{k}_{h}(s_{h}^{k})\}_{h=1,k=0}^{H-1,K-1}$ defined as
$\tilde{\eta}^{k}_{h}(s_{h}^{k})=\tilde{\Delta}^{k}_{h}(s_{h}^{k})-\mathbb{E}\left[\tilde{\Delta}^{k}_{h}(s_{h}^{k})\big|\mathcal{F}_{h-1}^{k}\right],$

### v3

@@ 4.1 Model-based tail bound: UCBVI algorithm
4.2 Model-free tail bounds: optimistic $Q$-learning
We analyze the tail behavior of an optimistic $Q$-learning algorithm studied in (Jin et al., 2018; Yang et al., 2021).
We consider the $K$-dependent (KD) settings, with bonus functions defined as
$b_{h}(n)=4H^{1.5}\sqrt{\dfrac{\ln(2SAHK)+\mu K^{\alpha}}{n}}$
Here $n\geq 1$ is the visit count of the associated state–action pair, and $\alpha\in[0,1]$ and $\mu>0$ are tuning parameters.
In contrast to the UCBVI algorithm, the update equations in optimistic $Q$-learning depend on the choice of step sizes. We define the step-size sequence and its unrolled weights as
$\beta^{(t)}\;\coloneqq\;\frac{H+1}{H+t+1},\qquad\beta^{t\rightarrow t^{\prime}}\;\coloneqq\;\beta^{(t)}\prod_{j=t+1}^{t^{\prime}}(1-\beta^{(j)}),\qquad t\geq 0.$ (4)
Notice that $\sum_{t=0}^{t^{\prime}}\beta^{t\to t^{\prime}}=1$, with the convention that $\prod_{j=t^{\prime}+1}^{t^{\prime}}\equiv 1$.
Algorithm 2 Optimistic Q-learning
1:Input: $H,\mathcal{S},\mathcal{A}$, rewards $r_{h}:\mathcal{S}\times\mathcal{A}\to[0,1]$, initial $s_{0}$, bonus schedules $\{b_{h}(\cdot)\}$, step size schedule $\{\beta^{(t)}\}$.
2:Init: For all $(s,a,h)$, set $Q_{h}^{0}(s,a)=H$ and $V_{h}^{0}(s)=H$ and $n_{h}^{0}(s,a)=0$.
3:for $k=0,1,2,\dots,K-1$ do
4:  Planning: For all $(s,h)$:
$\pi_{h}^{k}(s)\in\arg\max_{a\in\mathcal{A}}Q_{h}^{k}(s,a).$
6:  Update: Set $V_{H}^{k}\equiv 0$. For all $h$:
$t=n_{h}^{k+1}(s_{h}^{k},a_{h}^{k})=n_{h}^{k}(s_{h}^{k},a_{h}^{k})+1$
$Q_{h}^{k+1}(s_{h}^{k},a_{h}^{k})\leftarrow(1-\beta^{(t)})\,Q_{h}^{k}(s_{h}^{k},a_{h}^{k})+\beta^{(t)}\Big(r_{h}(s_{h}^{k},a_{h}^{k})+V_{h+1}^{k}(s_{h+1}^{k})+b_{h}(t)\Big),$
$V_{h}^{k+1}(s_{h}^{k})\leftarrow\min\{H,\max_{a\in\mathcal{A}}Q_{h}^{k+1}(s_{h}^{k},a)\}.$
7:end for
Algorithm 2 maintains optimistic action-value estimates $\{Q_{h}^{k}(s,a)\}$ and state-value estimates $\{V_{h}^{k}(s)\}$ for each stage $h\in\{0,1,\dots,H-1\}$, together with visitation counts $n_{h}^{k}(s,a)$. It is initialized optimistically as $Q_{h}^{0}(s,a)=H$, $V_{h}^{0}(s)=H$ for all $(s,a,h)$, and $n_{h}^{0}(s,a)=0$.
At the start of episode $k$, the planning step forms a greedy policy with respect to the current $Q$-values. The execution step then rolls out $\pi^{k}=\{\pi_{h}^{k}\}_{h=0}^{H-1}$ from the fixed initial state $s_{0}$ in the unknown MDP with transitions $P_{h}^{*}$, generating a trajectory $(s_{h}^{k},a_{h}^{k},s_{h+1}^{k})_{h=0}^{H-1}$. After observing the trajectory, the update step performs on-trajectory temporal-difference updates. First, the algorithm sets the terminal value $V_{H}^{k}\equiv 0$. Then for each stage $h=0,1,\dots,H-1$, it increments the visit count of the experienced pair $(s_{h}^{k},a_{h}^{k})$, and updates only the corresponding $Q$-entry using step size $\beta^{(t)}\in(0,1]$ toward a bonus-augmented TD target:
$Q_{h}^{k+1}(s_{h}^{k},a_{h}^{k})=(1-\beta^{(t)})\,Q_{h}^{k}(s_{h}^{k},a_{h}^{k})+\beta^{(t)}\Big(r_{h}(s_{h}^{k},a_{h}^{k})+V_{h+1}^{k}(s_{h+1}^{k})+b_{h}\big(t\big)\Big),$
where $t=n_{h}^{k+1}(s_{h}^{k},a_{h}^{k})$ is the visitation count to state-action $(s,a)$ at horizon $h$ at the end of episode $k$. All other action-values are left unchanged. Finally, the state-value at the visited states are updated to the greedy value and are clipped to the horizon, i.e., $V_{h}^{k+1}(s_{h}^{k})=\min\{H,\max_{a\in\mathcal{A}}Q_{h}^{k+1}(s_{h}^{k},a)\}$, while $V_{h}^{k+1}(s)=V_{h}^{k}(s)$ for all other states. The optimism is induced by the initial over-estimates and the nonnegative bonus $b_{h}(\cdot)$, decreasing in the visit count to encourage exploration early and vanish as sampling increases.
The episode index $k$ in $Q_{h}^{k}$ and $V_{h}^{k}$ is used only for notational clarity (e.g., in the analysis) to emphasize the evolution of the estimates across episodes. In an actual implementation, it suffices to store and update a single table for $\{Q_{h}(s,a)\}$ and a single table for $\{V_{h}(s)\}$, updating their entries in-place as data are collected.
To mirror Theorem 1 in our model-based section, we state a parameterized tail bound indexed by $\alpha\in[0,1]$, with an explicit sub-Gaussian component plus a residual “bad event” term.
Theorem 2.
Fix $\alpha\in[0,1]$ and $\mu>0$. The tail of the cumulative regret $R_{K}$ of Algorithm 2 satisfies
$\mathbb{P}\!\Big(R_{K}\geq x\Big)\;\leq\;\exp\!\left(-\frac{\big(x-m_{K}^{\mathrm{mf}}\big)_{+}^{2}}{H^{3}K}\right)+\delta_{K}^{\mathrm{mf}},\qquad\forall x\geq 0,$ (5)
where the baseline level is
$m_{K}^{\mathrm{mf}}\;\coloneqq\;\frac{3242eSAH^{6}(\ln(2SAHK)+\mu K^{\alpha})}{\mathrm{gap}^{*}},$ (6)
and the residual term is
$\delta_{K}^{\mathrm{mf}}\;\coloneqq\;\exp\!\big(-\mu K^{\alpha}\big)$ (7)
Similar to Theorem 1, Theorem 2 decomposes tail of regret into a sub-Gaussian part plus a residual $\delta_{K}^{\mathrm{mf}}$. Finally, we derive a bound for the expected regret.
Proposition 2.
Under the same setting as in Theorem 2, we have
$\mathbb{E}[R_{K}]\ \leq\ m_{K}^{\mathrm{mf}}+2KH^{2}\delta_{K}^{\mathrm{mf}}$
where $m_{K}^{\mathrm{mf}}$ is given in (6) and $\delta_{K}^{\mathrm{mf}}$ in (7).
5 Proof sketch
We outline the proof of Theorem 1 and highlight where each term in the bound arises. The proof of Theorem 2 follows similar ideas and is presented in the Appendix.
Burn-in decomposition.
Split the cumulative regret into an initial “burn-in” part and the remainder:
$R_{K}=\underbrace{\sum_{k=0}^{\lceil K^{\gamma}\rceil-1}\!\Big(V_{0}^{*}(s_{0})-V_{0}^{\bm{\pi}^{k}}(s_{0})\Big)}_{R_{K}^{(0)}}+\underbrace{\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\!\Big(V_{0}^{*}(s_{0})-V_{0}^{\bm{\pi}^{k}}(s_{0})\Big)}_{R_{K}^{(1)}}.$
Since each episode regret is at most $H$, we have $R_{K}^{(0)}\leq H\lceil K^{\gamma}\rceil,$ which yields the additive shift $H\lceil K^{\gamma}\rceil$ in the final tail bound. This initial burn-in term is essential for the analysis of the horizon-independent algorithm. By tuning $\gamma$, we trade off $R_{K}^{(0)}$ and $R_{K}^{(1)}$, so that the order matches.
Bellman-style telescoping and a martingale decomposition.
Fix an episode $k\geq\lceil K^{\gamma}\rceil$ and define the value difference $\Delta_{h}^{k}(s)\coloneqq V_{h}^{*}(s)-V_{h}^{\bm{\pi}^{k}}(s)$. Along the realized trajectory $(s_{h}^{k},a_{h}^{k})_{h=0}^{H-1}$ one can write
$\Delta_{h}^{k}(s_{h}^{k})=\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})+\mathbb{E}\!\left[\Delta_{h+1}^{k}(s_{h+1}^{k})\mid\mathcal{F}_{h}^{k}\right],$
and define the martingale difference term
$M_{h}^{k}(s_{h}^{k})\;\coloneqq\;\Delta_{h}^{k}(s_{h}^{k})-\mathbb{E}\!\left[\Delta_{h}^{k}(s_{h}^{k})\mid\mathcal{F}_{h-1}^{k}\right].$
Summing over $h=0,\dots,H-1$ telescopes (using $\Delta_{H}^{k}\equiv 0$) to give
$V_{0}^{*}(s_{0})-V_{0}^{\bm{\pi}^{k}}(s_{0})=\sum_{h=0}^{H-1}\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})-\sum_{h=0}^{H-1}M_{h+1}^{k}(s_{h+1}^{k}).$ (8)
We introduce the $\mathrm{gap}$ function because it enables instance-dependent guarantees.
Clipping and the role of the global gap.
We define the clipping function
$\mathrm{clip}(x)=0\ \text{if }x<\mathrm{gap}^{*},\qquad\mathrm{clip}(x)=x\ \text{otherwise}.$
By definition of the global gap $\mathrm{gap}^{*}$, for every $(s,h)$ we have $\mathrm{gap}_{h}(s,a)\in\{0\}\cup[\mathrm{gap}^{*},\infty)$, hence $\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})=\mathrm{clip}(\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k}))$. Thus (8) becomes
$V_{0}^{*}(s_{0})-V_{0}^{\bm{\pi}^{k}}(s_{0})=\sum_{h=0}^{H-1}\mathrm{clip}\!\big(\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})\big)-\sum_{h=0}^{H-1}M_{h+1}^{k}(s_{h+1}^{k}).$
The purpose of $\mathrm{clip}(\cdot)$ is that it lets us upper bound only the suboptimal-action contributions via bonuses, while treating the remaining fluctuations as martingale noise.
A high-probability “good event” implying optimism.
Define the per-episode good event
$\mathcal{G}^{k}\;=\;\bigcap_{h=0}^{H-2}\bigcap_{s,a}\left\{\|P_{h}^{*}(\cdot\mid s,a)-\hat{P}_{h}^{k}(\cdot\mid s,a)\|_{1}\leq\frac{2b_{h}^{k}(n_{h}^{k}(s,a))}{H-h-1}\right\}.$
On $\mathcal{G}^{k}$, a standard backward induction yields optimism:
$Q_{h}^{k}(s,a)\geq Q_{h}^{*}(s,a),\qquad V_{h}^{k}(s)\geq V_{h}^{*}(s),\qquad\forall(s,a,h).$ (9)
Let $\bar{\mathcal{G}}^{K}\coloneqq\cap_{k=\lceil K^{\gamma}\rceil}^{K-1}\mathcal{G}^{k}$ denote the good event over all post burn-in episodes.
Bounding the bad-event probability.
Using a union bound over $(s,a,h)$ and over visit counts $n\leq K$, and applying (Weissman et al., 2003, Theorem 2.1) to control $\|\hat{P}_{h}^{k}(\cdot\mid s,a)-P_{h}^{*}(\cdot\mid s,a)\|_{1}$ gives
$\mathbb{P}\big((\bar{\mathcal{G}}^{K})^{c}\big)\leq SAHK\cdot\begin{cases}\exp(-2\mu K^{\alpha}),&\texttt{KD},\\ \exp(-2\mu K^{\gamma\alpha}),&\texttt{KI},\end{cases}\;=\;\delta_{K}^{\mathrm{mb}}(\gamma).$
Regret on the good event $\leq$ deterministic “bonus budget” + martingale.
On $\bar{\mathcal{G}}^{K}$, optimism (9) lets us compare the clipped gap to the algorithmic optimism gap $\tilde{\Delta}_{h}^{k}(s)\coloneqq V_{h}^{k}(s)-V_{h}^{\bm{\pi}^{k}}(s)$. More specifically, on $\bar{\mathcal{G}}^{K}$, $\mathrm{clip}(\mathrm{gap}(s,a))\leq\mathrm{clip}(\tilde{\Delta}_{h}^{k}(s))$. A one-step recursion for $\tilde{\Delta}_{h}^{k}(s_{h}^{k})$ implies
$\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}\;\leq\;2\,b_{h}^{k}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))\mathbbm{1}\{\mathcal{G}^{k}\}+\mathbb{E}\!\left[\tilde{\Delta}_{h+1}^{k}(s_{h+1}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}\mid\mathcal{F}_{h}^{k}\right],$
and telescoping over $h$ yields an upper bound of the form
$\tilde{\Delta}_{h}^{k}(s_{h}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}\;\lesssim\;\sum_{l=h}^{H-1}b_{l}^{k}(n_{l}^{k}(s_{l}^{k},a_{l}^{k}))\mathbbm{1}\{\mathcal{G}^{k}\}\;-\;(\text{martingale terms}).$
Combining this with the clipping inequality $\mathrm{clip}\!\left(\sum_{i=1}^{H}a_{i}\right)\leq\sum_{i=1}^{H}\mathrm{clip}(Ha_{i})$ for all $a_{i}\geq 0$, one obtains a bound of the form
$R_{K}^{(1)}\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}\;\leq\;\underbrace{\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}(h+1)\,\mathrm{clip}\!\Big(4H\,b_{h}^{k}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))\Big)}_{\text{bonus budget}}\;-\;\underbrace{\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\bar{M}_{h+1}^{k}}_{\text{martingale}},$ (10)
where $\{\bar{M}_{h+1}^{k}\}$ is a martingale difference sequence.
Bounding the bonus budget by $m_{K}^{\mathrm{mb}}$ using the gap.
Because $\mathrm{clip}(x)=0$ for $x<\mathrm{gap}^{*}$, the summand in the bonus budget is nonzero only while
$4H\,b_{h}^{k}(n)\;\geq\;\mathrm{gap}^{*},\quad\text{i.e.,}\quad n\;\leq\;\frac{16H^{4}(0.5S\ln 2+\mu K^{\alpha})}{{\mathrm{gap}^{*}}^{2}}.$
Thus, for each $(s,a,h)$, only the first $\bar{n}=\lfloor 16H^{4}(0.5S\ln 2+\mu K^{\alpha})/{\mathrm{gap}^{*}}^{2}\rfloor$ visits contribute, and summing $\sum_{n=1}^{\bar{n}}\frac{1}{\sqrt{n}}\leq 2\sqrt{\bar{n}}$ gives
$\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}(h+1)\,\mathrm{clip}\!\Big(4H\,b_{h}^{k}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))\Big)\;\leq\;\frac{32H^{6}SA(0.5S\ln 2+\mu K^{\alpha})}{\mathrm{gap}^{*}}\;=\;m_{K}^{\mathrm{mb}}.$
Sub-Gaussian tail from Azuma–Hoeffding.
The martingale increments $\bar{M}_{h+1}^{k}$ are uniformly bounded by $5H^{2}$. There are at most $HK$ such increments, hence Azuma–Hoeffding yields
$\mathbb{P}\!\left(-\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\sum_{h=0}^{H-1}\bar{M}_{h+1}^{k}\geq t\right)\leq\exp\!\left(-\frac{t^{2}}{50H^{5}K}\right).$
Setting $t=(x-H\lceil K^{\gamma}\rceil-m_{K}^{\mathrm{mb}})_{+}$ and using (10) gives the sub-Gaussian term in the theorem.
Combine with the bad event.
Finally, decompose
$\mathbb{P}(R_{K}\geq x)\leq\mathbb{P}\big((\bar{\mathcal{G}}^{K})^{c}\big)+\mathbb{P}\big(\{R_{K}\geq x\}\cap\bar{\mathcal{G}}^{K}\big),$
and combining with the above inequalities we get the result in Theorem 1.
@@ 6.1 Proof of Theorem 1
Recall that $\bm{P}^{*}=\{P^{*}_{h}\}_{h=0}^{H-2}$ denotes the true transition kernel. At episode $k$, the algorithm computes a greedy policy $\pi^{k}=\{\pi^{k}_{h}\}_{h=0}^{H-1}$ using an optimistic estimate of the optimum value function $\{V^{*}_{h}\}_{h=0}^{H-1}$ based on data collected up to episode $k$. Recall that $R_{K}$ denote the cumulative regret incurred up to episode K1K-1. Then, writing $s_{0}^{k}\equiv s_{0}$, we have:
@@
$=\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})+\Delta^{k}_{h+1}(s_{h+1}^{k})-M^{k}_{h+1}(s_{h+1}^{k}),$
where the sequence $\{M^{k}_{h}(s_{h}^{k})\}_{h=1,k=0}^{H-1,K-1}$ defined as
$M^{k}_{h}(s_{h}^{k})=\Delta^{k}_{h}(s_{h}^{k})-\mathbb{E}\left[\Delta^{k}_{h}(s_{h}^{k})\big|\mathcal{F}_{h-1}^{k}\right]$
is a martingale difference sequence with respect to the filtration $\{\mathcal{F}_{h-1}^{k}\}_{h=1,k=0}^{H-1,K-1}$. For the sake of notational simplicity, we define $M^{k}_{H}\equiv 0$ for all $k\in[K]$.
Let $\upsilon^{k}\in\{0,1,\cdots,H-1,H\}$ denote the first horizon at which a non-optimal action is taken at episode $k$, i.e.,
$\upsilon^{k}\coloneqq\inf\{h:a_{h}^{k}\notin\mathcal{A}^{*}_{h}(s_{h}^{k})\}$
with the convention that $\inf\emptyset=H$. Notice that for any $h<\upsilon^{k}$, we have $\mathrm{gap}_{h}(s_{h}^{k},a_{h}^{k})=0$. Using a telescopic sum, noting that $\Delta^{k}_{H}\equiv 0$, we obtain
$\Delta^{k}_{0}(s_{0}^{k})$ $=\Delta^{k}_{\upsilon^{k}}(s_{\upsilon^{k}}^{k})-\sum_{h=0}^{{\upsilon^{k}}-1}M^{k}_{h+1}(s_{h+1}^{k})$ (11)
Note that $\Delta^{k}_{\upsilon^{k}}(s_{\upsilon^{k}}^{k})=V_{\upsilon^{k}}^{*}(s_{\upsilon^{k}}^{k})-V_{\upsilon^{k}}^{k}(s_{\upsilon^{k}}^{k})+V_{\upsilon^{k}}^{k}(s_{\upsilon^{k}}^{k})-V_{\upsilon^{k}}^{\bm{\pi}^{k}}(s_{\upsilon^{k}}^{k})$. The bonus function $b_{h}^{k}(\cdot)$ is chosen so that $V_{h}^{k}$ provides an optimistic estimate of $V_{h}^{*}$, i.e. $V_{h}^{k}\geq V_{h}^{*}$, as long as the estimated transition kernel $\hat{P}_{h^{\prime}}^{k}$ and the true transition kernel $P_{h^{\prime}}^{*}$ remain close for all horizons $h^{\prime}\geq h$. This combined with the monotonicity of the $\mathrm{clip}(\cdot)$ function provides an upper bound for the regret at episode $k\geq\lceil K^{\gamma}\rceil$. We define the “good event at episode kk” as
@@
We proceed with bounding the cumulative regret $R_{K}^{(1)}=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left(V_{0}^{*}(s_{0}^{k})-V_{0}^{\bm{\pi}_{k}}(s_{0}^{k})\right)$, under the event $\bar{\mathcal{G}}^{K}$.
$R_{K}^{(1)}\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ $=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left[\Delta^{k}_{\upsilon^{k}}(s_{\upsilon^{k}}^{k})-\sum_{h=0}^{{\upsilon^{k}}-1}M^{k}_{h+1}(s_{h+1}^{k})\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ (by (11))
$=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left[V_{\upsilon^{k}}^{*}(s_{\upsilon^{k}}^{k})\pm V_{\upsilon^{k}}^{k}(s_{\upsilon^{k}}^{k})-V_{\upsilon^{k}}^{\bm{\pi}^{k}}(s_{\upsilon^{k}}^{k})-\sum_{h=0}^{{\upsilon^{k}}-1}M^{k}_{h+1}(s_{h+1}^{k})\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$
$=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left[\left(V_{\upsilon^{k}}^{*}(s_{\upsilon^{k}}^{k})\pm V_{\upsilon^{k}}^{k}(s_{\upsilon^{k}}^{k})-V_{\upsilon^{k}}^{\bm{\pi}^{k}}(s_{\upsilon^{k}}^{k})\right)\mathbbm{1}\{\mathcal{G}^{k}\}-\sum_{h=0}^{{\upsilon^{k}}-1}M^{k}_{h+1}(s_{h+1}^{k})\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$
$\leq\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left[\left(V_{\upsilon^{k}}^{k}(s_{\upsilon^{k}}^{k})-V_{\upsilon^{k}}^{\bm{\pi}^{k}}(s_{\upsilon^{k}}^{k})\right)\mathbbm{1}\{\mathcal{G}^{k}\}-\sum_{h=0}^{{\upsilon^{k}}-1}M^{k}_{h+1}(s_{h+1}^{k})\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\}$ (by (12))
$=\sum_{k=\lceil K^{\gamma}\rceil}^{K-1}\left[\tilde{\Delta}_{\upsilon^{k}}^{k}(s_{\upsilon^{k}}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}-\sum_{h=0}^{{\upsilon^{k}}-1}M^{k}_{h+1}(s_{h+1}^{k})\right]\mathbbm{1}\{\bar{\mathcal{G}}^{K}\},$ (13)
@@
$=2\mathbbm{1}\{\mathcal{G}^{k}\}b^{k}_{h}(n_{h}^{k}(s_{h}^{k},a_{h}^{k}))+\tilde{\Delta}_{h+1}^{k}(s_{h+1}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}-\tilde{M}_{h+1}^{k}(s_{h+1}^{k})\mathbbm{1}\{\mathcal{G}^{k}\}.$ (14)
@@
where $(a)$ follows from Hölder’s inequality, and noting that $\sum_{s}(P_{h}^{*}(s\mid s_{h}^{k},a_{h}^{k})-\hat{P}_{h}^{k}(s\mid s_{h}^{k},a_{h}^{k}))=0$, and the sequence $\{\tilde{M}^{k}_{h}(s_{h}^{k})\}_{h=1,k=0}^{H-1,K-1}$ defined as
$\tilde{M}^{k}_{h}(s_{h}^{k})=\tilde{\Delta}^{k}_{h}(s_{h}^{k})-\mathbb{E}\left[\tilde{\Delta}^{k}_{h}(s_{h}^{k})\big|\mathcal{F}_{h-1}^{k}\right],$